# Gravitational and electromagnetic Cherenkov radiation constraints in modified dispersion relations

### v1

@@ 1 Introduction
Electromagnetic Cherenkov radiation occurs when an electrically charged particle travelling through an optical medium has a velocity $v$ larger than the phase velocity of light in that medium $c_{\gamma}$. This phenomenon was first detected by Cherenkov in 1934 Cherenkov_1934 , and a theoretical explanation was given by Frank and Tamm three years later Tamm_1937 . Cherenkov radiation (CR) can be understood microscopically in the following way Schreck_2018 . When the particle travels at a velocity $v\leq c_{\gamma}$, the polarized atoms close to the trajectory emit out of phase radiation, leading to destructive interference. However, when the particle has a velocity $v>c_{\gamma}$, the wave trains emitted by atoms and molecules are in phase, and the constructive interference taking place in a Mach cone, known as the Cherenkov cone, results in the coherent radiation observed in the direction perpendicular to the cone. Likewise, this phenomenon can be interpreted both qualitatively and quantitatively as an emission process $a\to a+\gamma$ where a charged particle $a$ emits a photon $\gamma$: the particle outruns the electromagnetic field, causing the emission of radiation because of the accumulation of wavefronts propagating from the particle Schreck_2018 ; Moore_2001 . The rate of energy loss in an optical medium with refractive index $n$ is described classically by the Frank-Tamm formula Tamm_1937 , and receives small corrections when considering quantum effects cox1944momentum . Moreover, this formula can also be obtained from the quantum field theory formalism Moore_2001 . Anyway, for the process to be allowed the velocity of the particle must be greater than the phase velocity of light and, therefore, Cherenkov emission is not possible in vacuum in the framework of special relativity (SR) since $c_{\gamma}=c$.
@@
Nevertheless, over the last decades there has been a growing interest in studying theoretical frameworks which suggest that Lorentz invariance (LI) may not be an exact symmetry of nature at all energy scales Mattingly_2005 . Such theories appear mainly in the context of quantum gravity (QG) Addazi_2022 ; batista2023white where, if LI is violated, the Planck energy $E_{\text{Pl}}\approx 10^{19}\,\text{GeV}$ is expected to be the scale where this symmetry is strongly violated. Although there is a large energy gap between the highest energy particles detected, those are Ultra High Energy Cosmic Rays (UHECRs) $\approx 10^{11}\,\text{GeV}$, and the Planck scale, there should be an interpolation of LI violation (LIV) to the low energy regime, where these particles could be sensitive to small departures from LI. Focusing on astrophysical phenomena, the most straightforward way to implement LIV is to consider modified dispersion relations (MDRs) for particles maintaining the $4-$momentum conservation laws. In this scenario, kinematics of a wide variety of processes may be affected in what are called threshold effects depending on the MDRs considered. These effects could entail the shift of the energy thresholds or add new thresholds to processes already predicted in the LI scenario, or even allow completely new reactions.
@@
Vacuum CR may be allowed in a LI violating frame, and it has been exhaustively studied in the electromagnetic sector Coleman_1997 ; Coleman_1999 ; Jacobson_2003 ; Jacobson_2006 ; Gagnon_2004 to impose restrictive constraints to LI violating parameters (LIVPs) using high energy astrophysical particles. This idea has also been extended to study the vacuum gravitational CR Moore_2001 ; Elliott_2005 ; Kimura_2012 ; Kiyota_2015 ; Kosteleck_2015 , where an arbitrary particle can lose energy due to the emission of gravitons rather than photons. With the recent observations of gravitational waves (GWs) Yunes_2016 ; Abbott_2017 , combined constraints for LIVPs were obtained in the gravitational sector. These studies, however, only modify the dispersion relation in the gravitational sector or the matter sector, but combined effects have not been investigated yet with non-trivial MDRs. The main aim of the present paper is to consider MDRs both in the gravitational and the matter sector simultaneously to obtain combined restrictions in the LIVPs taking into account the existence or absence of vacuum gravitational CR.
@@
The remainder of this work is structured as follows. In section 2 we propose a phenomenological MDR and study the kinematics of the generalised Cherenkov radiation (GCR) process $a\to a+b$. In section 3 we apply the results obtained in section 2 to compute the rate of energy loss of the particle (section 3.1) for vacuum electromagnetic and gravitational CR (sections 3.2 and 3.3). Section 4 is devoted to impose constraints in LIVPs using UHECRs, and, finally, in section 5 we discuss the main results obtained and its limitations. Along this work we use natural units, i.e., we set $\hbar=c=1$.
@@ 2.1 General aspects of the modified dispersion relation
Nowadays it is not yet clear how LI might be broken, if it is the case. However, there are hints about the kind of phenomenology it could produce. An interesting example is the possibility of new particle decays due to LIV. In order to study the kinematics of these processes, it has previously been proposed Coleman_1999 that each particle $a$ has, besides its own mass $m_{a}$, its own maximum velocity $c_{a}$, which is asymptotically achievable if the particle has a non-vanishing mass. The dispersion relation in this scenario is given by
@@
where $E_{a}$ and $\bm{p}_{a}$ are the energy and 3-momentum of the particle, respectively. On the other hand, the Cherenkov effect is usually studied in terms of the refractive index $n$ of an optical medium, which modifies the maximum attainable velocity of light in vacuum $c$ as $c_{\gamma}=cn^{-1}$. Hereafter we will set $c=1$. Following this idea, we consider that it is convenient to parameterize the deviation of the maximum attainable velocity $c_{a}$ for each particle species in terms of a particular refractive index $n_{a}$, so that $c_{a}=n_{a}^{-1}$ and equation (1) can be written as
@@
where there is no summation in $\alpha$ as we take into account only the dominant term in the deviation from unity. Both the coefficient $\mathbb{A}^{(\alpha)}$, with units $[\mathbb{A}^{(\alpha)}]=[E^{-\alpha}]$, and the exponent $\alpha$ are allowed to be different for each particle species. From now on we shall denote $\mathbb{A}\equiv\mathbb{A}^{(\alpha)}$ except when discussing different values of $\alpha$. As LI is almost an exact symmetry we will assume that $|\mathbb{A}E_{\bm{p}_{a}}^{\alpha}|\ll 1$. Replacing the refractive index (3) in the MDR (2) and considering terms up to first order in $|\mathbb{A}E_{\bm{p}_{a}}^{\alpha}|$ we get
@@
Even though $m_{a}$ and $|\mathbb{A}E_{\bm{p}_{a}}^{\alpha}|$ are independent scales, we will see that it is safe to expand only in the LI violating term for high energy particles, hence no multiscale problems arise from this analysis. Note that equation (4) is invariant under translations and rotations but not under Lorentz boosts due to the dependence of $n_{a}$ on the energy of the particle. Thus, this MDR only holds in a particular preferred frame, which is usually chosen to be in rest with respect to the CMB. The relative velocity between Earth and the CMB is of order $\mathcal{O}(10^{-3})$ Planck_2018 and its effects can be neglected when imposing constraints on the LIVPs.
It should be noted that for the MDR given by equation (2) the group velocity and the phase velocity are different in general. The maximum attainable velocity of the particle $c_{a}=n_{a}^{-1}=1-\mathbb{A}p^{\alpha}+\mathcal{O}(\mathbb{A}^{2})$ is referred to the phase velocity, which differs from the group velocity defined as $v_{a}(E)=\partial E/\partial p$. Indeed, for high energy particles with $p\gg m$ one has, using equation (4) and neglecting the mass in the last term,
@@
In this work we are interested in studying the existence of entirely new reactions that are forbidden in SR; in particular we will consider the GCR process (see figure 1), i.e., the two body decay process $a\to a+b$ where a particle $a$ emits a massless particle $b$ [^1] Here we mention that forbidden processes in the LI scenario cannot be studied in the scope of Doubly Special Relativity (DSR). This is because the deformed $4-$momentum composition produces cancellations with the LI violating terms in such a way that this phenomenology is not possible Addazi_2022 ..
Figure 1: Tree level GCR process where a particle $a$ (continuous line) emits a massless particle $b$ (wiggly line).
@@ 2.2 Derivation of the Cherenkov angle
One of the key ingredients when studying CR is the angle of emission $\theta_{c}$ between the $3-$momentum $\bm{p}$ of the initial particle $a$ and the $3-$momentum $\bm{k}$ of the emitted particle $b$. If we define $E_{\bm{p}}:=E_{a}(\bm{p})$, $E_{\bm{p^{\prime}}}:=E_{a}(\bm{p}^{\prime})$ and $E_{\bm{k}}:=E_{b}(\bm{k})$, where $\bm{p}^{\prime}$ is the $3-$momentum of the particle $a$ after the emission (see figure 1), then the conservation of energy and $3-$momentum, which are written as
@@
allow us to obtain $\theta_{c}$ in terms of the module of the $3$-momenta $\bm{p}$ and $\bm{k}$. Note that equation (7) assumes the usual $4-$momentum composition, hence DSR theories are excluded from this analysis.
Let us now use the MDR given by equation (2) with refractive index (3) to obtain the Cherenkov angle $\theta_{c}$ or, equivalently, $\cos{\theta_{c}}$ ). Let $m$ be the mass of particle $a$ (we remind the reader that $b$ is massless) and denote $n_{\bm{p}}:=n_{a}(\bm{p})$, $n_{\bm{p}-\bm{k}}:=n_{a}(\bm{p}-\bm{k})$ and $n_{\bm{k}}:=n_{b}(\bm{k})$, where we already applied the conservation of $3-$momentum. From equation (2), we see that the energies of the particles in the GCR process are given by
@@
where we have used $(\bm{p}-\bm{k})^{2}=p^{2}+k^{2}-2pk\cos{\theta_{c}}$ ) and denoted the module of the $3-$momentum of the particles $a$ and $b$ as $p$ and $k$, respectively. On the other hand, combining the conservation of energy and $3-$momentum we have $E_{\bm{p}-\bm{k}}^{2}=[E_{\bm{p}}-E_{\bm{k}}]^{2}$. Substituting here the energies given by equation (8), we get
@@
Note that in the LI scenario, which corresponds to $n_{a}=n_{b}=1$, the RHS of equation (9) is always greater than the LHS unless $\cos{\theta_{c}}>1$; hence, the process is forbidden. The limit case where the particle $a$ has no mass is of no interest when $n_{a}=n_{b}=1$, as the emission rate, which we compute in section 3, vanishes for $\cos{\theta_{c}}=1$.
@@
We remark that for the massless particle we are ignoring a small mass term $m_{b}$, and terms that induce birefringence or attenuation, which could be present in alternative theories of gravity for the graviton Cembranos_2019 or in the Standard Model extension for gauge bosons. Assuming that $|\mathbb{A}E_{\bm{p}}^{\alpha}|,|\mathbb{A}E_{\bm{p}-\bm{k}}^{\alpha}|,|\mathbb{B}E_{\bm{k}}^{\alpha}|\ll 1$ we can replace these refractive indexes in equation (9) and expand up to first order to obtain
@@
This is an implicit equation in $\cos{\theta}$ ) as the energy $E_{\bm{p}-\bm{k}}$ depends on $(\bm{p}-\bm{k})^{2}$, which in turn depends on $\cos{\theta}$ ).
In section 4 we will consider UHECR observations to constrain the LI breaking parameters $\mathbb{A}$ and $\mathbb{B}$. For these particles $p\gg m$. In addition, it can be checked a posteriori that the threshold is greatly modified, that is, the momentum is much larger than the right hand side of equation (6). We can then consider terms up to $\mathcal{O}(m^{2}/p^{2})$ and neglect first order products of the LI violating terms with the quotient $m^{2}/p^{2}$. Applying these additional approximations in equation (11) we get
@@
As $\Theta_{c}$ is already of first order in the LI violating terms and $m^{2}/p^{2}$, then the $3-$vectors $\bm{p}$ and $\bm{k}$ involved in the energies must be considered collinear and no second order corrections are taken into account. In other words, $\Theta_{c}$ does not depend on $\cos{\theta_{c}}$ ) and, consequently, equation (12) completely characterizes the Cherenkov angle $\theta_{c}$ in terms of the initial $3-$momentum of the particle $a$ and the $3-$momentum carried off by the particle $b$.
@@ 2.3 Threshold condition
Attending to equation (12), the Cherenkov process $a\to a+b$ is allowed as long as $\Theta_{c}>0$ such that $\cos{\theta_{c}}<1$. Hence the threshold is determined by the equation
@@
The study of the threshold condition is easier to carry out if the energies in equation (13) are substituted by the momentum, which is correct up to first order. Thus, we obtain
@@
In what follows we will consider $k\leq p$, which is satisfied particularly when studying the threshold condition. Indeed, if $k>p$, then the massless particle would be anti-parallel to the incoming particle, but this is forbidden by the threshold theorem Mattingly_2003 : if $E_{\bm{p}}$ is a strictly monotonically increasing function of $p$ for $p>0$ for all particles, then all thresholds for processes with two particle final states occur when the final momentum are parallel. This theorem has to be satisfied in our case since, in the first place, as we have previously discussed, $E_{\bm{p}}$ is a rotational-invariant function of $\bm{p}$, so it is only function of the modulus of the 3-momentum $p$; in the second place, it is a strictly monotonically increasing function of $p$, as the LI violating terms are much smaller than the momentum of the particle.
@@
    $\mathbb{A}>0$ and $\mathbb{B}<0$. It is easy to see from equation (15) that $\Theta_{c}<0$ for all $k<p$, as the energy of the incoming particle is always greater than the energy of the particles in the final state. Hence, Cherenkov emission is forbidden. In this scenario $n_{a}>n_{b}$ and, therefore, $c_{a}<c_{b}$. So, the intuitive condition that the particle $a$ must have a greater phase velocity than the particle $b$ in order for the process to be allowed is never fulfilled.
@@
    $\mathbb{A}<0$ and $\mathbb{B}>0$. In this case we see that all the terms in equation (15) are positive except the correction $m^{2}/(2p^{2})$ due to the mass of the particle. For particles with energies many orders of magnitude greater than the one established by the RHS of equation (6), this term might be neglected and, thus, Cherenkov emission is allowed for $k\lesssim p$. Now we have that $n_{a}<n_{b}$ ($c_{a}>c_{b}$), which fits with the idea that the process is possible since the particle $a$ has a greater phase velocity than the particle $b$.
@@
Before doing so, we can study the relation between $\mathbb{A}$ and $\mathbb{B}$ when the emitted particle $b$ has arbitrarily low momentum, $k\to 0$, or carries off all of the momentum of the particle $a$, $k\to p$. In the former case, it is easy to see from equation (15) that Cherenkov emission is possible if:
@@
The threshold condition is recovered setting an equal sign in equations (17) and (18).
@@
It is easy to check that, for the physical momentum $k\in[0,p]$, $g^{\prime}(k)>0$ and $g(0)$, $g(p)>0$; hence $g(k)>0$. The remaining terms in equation (19) are positive, so we also have that $f^{\prime}(k)>0$; in other words, $f(k)$ is a monotonically increasing function. Therefore, if $f(0)<0$ and $f(p)>0$, the solution $k_{\text{th}}$ is unique. The condition $f(0)<0$ is satisfied when equation (17) does not hold, and $f(p)>0$ when equation (18) is fulfilled. The threshold momentum $k_{\text{th}}$ here establishes the minimum momentum that particle $b$ must have in order to have Cherenkov emission with momentum $k\in[k_{\text{th}},p]$, and it is permitted as long as $k_{\text{th}}<p$. The limit case $k_{\text{th}}=p$ occurs setting an equal sign in equation (18).
The prove is similar when $\mathbb{A}<0$ and $\mathbb{B}<0$. In this case one can check that $f(k)$ is a monotonically decreasing function, and so there exists a unique solution $k_{\text{th}}$ when $f(0)>0$ and $f(p)<0$. The first condition is met when equation (17) is satisfied, and the second one when equation (18) is not fulfilled. Hence, $k_{\text{th}}$ establishes the maximum momentum that particle $b$ can have up to where the Cherenkov effect is allowed, and thus $k\in[0,k_{\text{th}}]$. The process is permitted when $k_{\text{th}}>0$, and the limit case $k_{\text{th}}=0$ is given setting an equal sign in equation (17).
@@
Table 1: In this table we show the threshold momentum $k_{\text{th}}$ for the particle $b$ that satisfy the condition $f(k)=0$ together with the conditions that $\mathbb{A}$ and $\mathbb{B}$ should satisfy for the process to be allowed. The $\pm$ signs in the quadratic solutions depend on the signs of $\mathbb{A}$ and $\mathbb{B}$, and these are chosen so that $k_{\text{th}}$ is positive for those LIVPs that fulfill either the condition (17) or (18). We emphasize again that $k_{\text{th}}$ establishes the minimum momentum of the particle $b$ when $\mathbb{A},\mathbb{B}>0$, and a maximum momentum when $\mathbb{A},\mathbb{B}<0$.
The threshold condition $\Theta_{c}=0$ does not have a general solution $k_{\text{th}}$ for arbitrary $\alpha$ and $\beta$ since it is a polynomial equation of degree $\max\{\alpha+1,\beta+1\}$. Nevertheless, particular solutions for small values of $\alpha$ and $\beta$ can be derived, and further simplified in the massless limit $m\to 0$. For example, consider the case where $\alpha=1$ and $\beta=0$. One easily checks from equation (16) that the threshold momentum for $m=0$ is given by

### v2

@@ 1 Introduction
Electromagnetic Cherenkov radiation occurs when an electrically charged particle traveling through an optical medium has a velocity $v$ larger than the phase velocity of light in that medium $c_{\gamma}$. This phenomenon was first detected by Cherenkov in 1934 Cherenkov_1934 , and a theoretical explanation was given by Frank and Tamm three years later Tamm_1937 . Cherenkov radiation (CR) can be understood microscopically in the following way Schreck_2018 . When the particle travels at a velocity $v\leq c_{\gamma}$, the polarized atoms close to the trajectory emit out of phase radiation, leading to destructive interference. However, when the particle has a velocity $v>c_{\gamma}$, the wave trains emitted by atoms and molecules are in phase, and the constructive interference taking place in a Mach cone, known as the Cherenkov cone, results in the coherent radiation observed in the direction perpendicular to the cone. Likewise, this phenomenon can be interpreted both qualitatively and quantitatively as an emission process $a\to a+\gamma$ where a charged particle $a$ emits a photon $\gamma$: the particle outruns the electromagnetic field, causing the emission of radiation because of the accumulation of wavefronts propagating from the particle Schreck_2018 ; Moore_2001 . The rate of energy loss in an optical medium with refractive index $n$ is described classically by the Frank-Tamm formula Tamm_1937 , and receives small corrections when considering quantum effects cox1944momentum . Moreover, this formula can also be obtained from the quantum field theory formalism Moore_2001 . Anyway, for the process to be allowed the velocity of the particle must be greater than the phase velocity of light and, therefore, Cherenkov emission is not possible in vacuum in the framework of special relativity (SR) since $c_{\gamma}=c$.
@@
Nevertheless, over the last decades there has been a growing interest in studying theoretical frameworks which suggest that Lorentz invariance (LI) may not be an exact symmetry of nature at all energy scales Mattingly_2005 . Such theories appear mainly in the context of quantum gravity (QG) Addazi_2022 ; batista2023white where, if LI is violated, the Planck energy $E_{\text{Pl}}\approx 10^{19}\,\text{GeV}$ is expected to be the scale where this symmetry is strongly violated. Although there is a large energy gap between the highest energy particles detected, those are ultra high energy cosmic rays (UHECRs) $\approx 10^{11}\,\text{GeV}$, and with the Planck scale, there should be an interpolation of LI violation (LIV) to the low energy regime, where these particles could be sensitive to small departures from LI. Focusing on astrophysical phenomena, the most straightforward way to implement LIV is to consider modified dispersion relations (MDRs) for particles maintaining the 4-momentum conservation laws. In this scenario, kinematics of a wide variety of processes may be affected in what are called threshold effects depending on the MDRs considered. These effects could entail the shift of the energy thresholds or add new thresholds to processes already predicted in the LI scenario, or even allow completely new reactions.
@@
Vacuum CR may be allowed in a LI violating frame, and it has been exhaustively studied in the electromagnetic sector Coleman_1997 ; Coleman_1999 ; Jacobson_2003 ; Jacobson_2006 ; Gagnon_2004 to impose restrictive constraints to LI violating parameters (LIVPs) using high energy astrophysical particles. This idea has also been extended to study the vacuum gravitational CR Moore_2001 ; Elliott_2005 ; Kimura_2012 ; Kiyota_2015 ; Kosteleck_2015 , where an arbitrary particle can lose energy due to the emission of gravitons rather than photons. With the recent observations of gravitational waves (GWs) Yunes_2016 ; Abbott_2017 , combined constraints for LIVPs were obtained in the gravitational sector. These studies, however, only modify the dispersion relation in the gravitational sector or the matter sector, but combined effects have not been investigated yet with nontrivial MDRs. The main aim of the present paper is to consider MDRs both in the gravitational and the matter sector simultaneously to obtain combined restrictions in the LIVPs taking into account the existence or absence of vacuum gravitational CR.
@@
The remainder of this work is structured as follows. In Sec. 2 we propose a phenomenological MDR and study the kinematics of the generalized Cherenkov radiation (GCR) process $a\to a+b$. In Sec. 3 we apply the results obtained in Sec. 2 to compute the rate of energy loss of the particle (Sec. 3.1) for vacuum electromagnetic and gravitational CR (Secs. 3.2 and 3.3). Sec. 4 is devoted to impose constraints in LIVPs using UHECRs, and, finally, in Sec. 5 we discuss the main results obtained and its limitations. Along this work we use natural units, i.e., we set $\hbar=c=1$.
@@ 2.1 General aspects of the modified dispersion relation
Nowadays it is not yet clear how LI might be broken, if it is the case. However, there are hints about the kind of phenomenology it could produce. An interesting example is the possibility of new particle decays due to LIV. In order to study the kinematics of these processes, it has previously been proposed Coleman_1999 that each particle $a$ has, besides its own mass $m_{a}$, its own maximum velocity $c_{a}$, which is asymptotically achievable if the particle has a nonvanishing mass. The dispersion relation in this scenario is given by
@@
where $E_{a}$ and $\bm{p}_{a}$ are the energy and 3-momentum of the particle, respectively. On the other hand, the Cherenkov effect is usually studied in terms of the refractive index $n$ of an optical medium, which modifies the maximum attainable velocity of light in vacuum $c$ as $c_{\gamma}=cn^{-1}$. Hereafter we will set $c=1$. Following this idea, we consider that it is convenient to parametrize the deviation of the maximum attainable velocity $c_{a}$ for each particle species in terms of a particular refractive index $n_{a}$, so that $c_{a}=n_{a}^{-1}$ and Eq. (1) can be written as
@@
where there is no summation in $\alpha$ as we take into account only the dominant term in the deviation from unity. Both the constant coefficient $\mathbb{A}^{(\alpha)}$, with units $[\mathbb{A}^{(\alpha)}]=[E^{-\alpha}]$, and the exponent $\alpha$ are allowed to be different for each particle species. To illustrate the utility of Eq. (3), we point out that when $\alpha=0$, some MDRs predicted by different theoretical frameworks can be described, e.g. the setting characterized by isotropic coefficients in the photon sector found in Klinkhamer_2008 or in the fermion sector studied in Schreck_2017 . From now on we shall denote $\mathbb{A}\equiv\mathbb{A}^{(\alpha)}$ except when discussing different values of $\alpha$. As LI is almost an exact symmetry we will assume that $|\mathbb{A}E_{\bm{p}_{a}}^{\alpha}|\ll 1$. Replacing the refractive index (3) in the MDR (2) and considering terms up to first order in $|\mathbb{A}E_{\bm{p}_{a}}^{\alpha}|$ we get
@@
Even though $m_{a}$ and $|\mathbb{A}E_{\bm{p}_{a}}^{\alpha}|$ are independent scales, we will see that it is safe to expand only in the LI violating term for high energy particles, hence no multiscale problems arise from this analysis. Note that Eq. (4) is invariant under translations and rotations but not under Lorentz boosts due to the dependence of $n_{a}$ on the energy of the particle. Thus, this MDR only holds in a particular preferred frame, which is usually chosen to be at rest with respect to the CMB. The relative velocity between Earth and the CMB is of order $\mathcal{O}(10^{-3})$ Planck_2018 and its effects can be neglected when imposing constraints on the LIVPs.
It should be noted that for the MDR given by Eq. (2) the group velocity and the phase velocity are different in general. The maximum attainable velocity of the particle $c_{a}=n_{a}^{-1}=1-\mathbb{A}p^{\alpha}+\mathcal{O}(\mathbb{A}^{2})$ is referred to the phase velocity, which differs from the group velocity defined as $v_{a}(E)=\partial E/\partial p$. Indeed, for high energy particles with $p\gg m$ one has, using Eq. (4) and neglecting the mass in the last term,
@@
In this work we are interested in studying the existence of entirely new reactions that are forbidden in SR; in particular we will consider the GCR process (see Fig. 1), i.e., the two body decay process $a\to a+b$ where a particle $a$ emits a massless particle $b$ [^1] Here we mention that forbidden processes in the LI scenario cannot be studied in the scope of doubly special relativity (DSR). This is because the deformed 4-momentum composition produces cancellations with the LI violating terms in such a way that this phenomenology is not possible Addazi_2022 ..
Figure 1: Tree-level GCR process where a particle $a$ (continuous line) emits a massless particle $b$ (wiggly line).
@@ 2.2 Derivation of the Cherenkov angle
One of the key ingredients when studying CR is the angle of emission $\theta_{c}$ between the 3-momentum $\bm{p}$ of the initial particle $a$ and the 3-momentum $\bm{k}$ of the emitted particle $b$. If we define $E_{\bm{p}}:=E_{a}(\bm{p})$, $E_{\bm{p^{\prime}}}:=E_{a}(\bm{p}^{\prime})$ and $E_{\bm{k}}:=E_{b}(\bm{k})$, where $\bm{p}^{\prime}$ is the 3-momentum of the particle $a$ after the emission (see Fig. 1), then the conservation of energy and 3-momentum, which are written as
@@
allow us to obtain $\theta_{c}$ in terms of the module of the $3$-momenta $\bm{p}$ and $\bm{k}$. Note that Eq. (7) assumes the usual 4-momentum composition; hence, DSR theories are excluded from this analysis.
Let us now use the MDR given by Eq. (2) with refractive index (3) to obtain the Cherenkov angle $\theta_{c}$ or, equivalently, $\cos{\theta_{c}}$ ). Let $m$ be the mass of particle $a$ (we remind the reader that $b$ is massless) and denote $n_{\bm{p}}:=n_{a}(\bm{p})$, $n_{\bm{p}-\bm{k}}:=n_{a}(\bm{p}-\bm{k})$ and $n_{\bm{k}}:=n_{b}(\bm{k})$, where we already applied the conservation of 3-momentum. From Eq. (2), we see that the energies of the particles in the GCR process are given by
@@
where we have used $(\bm{p}-\bm{k})^{2}=p^{2}+k^{2}-2pk\cos{\theta_{c}}$ ) and denoted the module of the 3-momentum of the particles $a$ and $b$ as $p$ and $k$, respectively. On the other hand, combining the conservation of energy and 3-momentum we have $E_{\bm{p}-\bm{k}}^{2}=[E_{\bm{p}}-E_{\bm{k}}]^{2}$. Substituting here the energies given by Eq. (8), we get
@@
Note that in the LI scenario, which corresponds to $n_{a}=n_{b}=1$, the rhs of Eq. (9) is always greater than the lhs unless $\cos{\theta_{c}}>1$; hence, the process is forbidden. The limit case where the particle $a$ has no mass is of no interest when $n_{a}=n_{b}=1$, as the emission rate, which we compute in Sec. 3, vanishes for $\cos{\theta_{c}}=1$.
@@
We remark that for the massless particle we are ignoring a small mass term $m_{b}$, and terms that induce birefringence or attenuation, which could be present in alternative theories of gravity for the graviton Cembranos_2019 or in the Standard Model extension for gauge bosons. Assuming that $|\mathbb{A}E_{\bm{p}}^{\alpha}|,|\mathbb{A}E_{\bm{p}-\bm{k}}^{\alpha}|,|\mathbb{B}E_{\bm{k}}^{\alpha}|\ll 1$ we can replace these refractive indexes in Eq. (9) and expand up to first order to obtain
@@
This is an implicit Eq. in $\cos{\theta}$ ) as the energy $E_{\bm{p}-\bm{k}}$ depends on $(\bm{p}-\bm{k})^{2}$, which in turn depends on $\cos{\theta}$ ).
In Sec. 4 we will consider UHECR observations to constrain the LI breaking parameters $\mathbb{A}$ and $\mathbb{B}$. For these particles $p\gg m$. In addition, it can be checked a posteriori that the threshold is greatly modified, that is, the momentum is much larger than the right-hand side of Eq. (6). We can then consider terms up to $\mathcal{O}(m^{2}/p^{2})$ and neglect first order products of the LI violating terms with the quotient $m^{2}/p^{2}$. Applying these additional approximations in Eq. (11) we get
@@
As $\Theta_{c}$ is already of first order in the LI violating terms and $m^{2}/p^{2}$, then the three-vectors $\bm{p}$ and $\bm{k}$ involved in the energies must be considered collinear and no second order corrections are taken into account. In other words, $\Theta_{c}$ does not depend on $\cos{\theta_{c}}$ ) and, consequently, Eq. (12) completely characterizes the Cherenkov angle $\theta_{c}$ in terms of the initial 3-momentum of the particle $a$ and the 3-momentum carried off by the particle $b$.
@@ 2.3 Threshold condition
Attending to Eq. (12), the Cherenkov process $a\to a+b$ is allowed as long as $\Theta_{c}>0$ such that $\cos{\theta_{c}}<1$. Hence the threshold is determined by the equation
@@
The study of the threshold condition is easier to carry out if the energies in Eq. (13) are substituted by the momentum, which is correct up to first order. Thus, we obtain
@@
In what follows we will consider $k\leq p$, which is satisfied particularly when studying the threshold condition. Indeed, if $k>p$, then the massless particle would be anti-parallel to the incoming particle, but this is forbidden by the threshold theorem Mattingly_2003 : if $E_{\bm{p}}$ is a strictly monotonically increasing function of $p$ for $p>0$ for all particles, then all thresholds for processes with two particle final states occur when the final momentum are parallel. This theorem has to be satisfied in our case since, in the first place, as we have previously discussed, $E_{\bm{p}}$ is a rotational-invariant function of $\bm{p}$, so it is only a function of the modulus of the 3-momentum $p$; in the second place, it is a strictly monotonically increasing function of $p$, as the LI violating terms are much smaller than the momentum of the particle.
@@
    $\mathbb{A}>0$ and $\mathbb{B}<0$. It is easy to see from Eq. (15) that $\Theta_{c}<0$ for all $k<p$, as the energy of the incoming particle is always greater than the energy of the particles in the final state. Hence, Cherenkov emission is forbidden. In this scenario $n_{a}>n_{b}$ and, therefore, $c_{a}<c_{b}$. So, the intuitive condition that the particle $a$ must have a greater phase velocity than the particle $b$ in order for the process to be allowed is never fulfilled.
@@
    $\mathbb{A}<0$ and $\mathbb{B}>0$. In this case we see that all the terms in Eq. (15) are positive except the correction $m^{2}/(2p^{2})$ due to the mass of the particle. For particles with energies many orders of magnitude greater than the one established by the rhs of Eq. (6), this term might be neglected and, thus, Cherenkov emission is allowed for $k\lesssim p$. Now we have that $n_{a}<n_{b}$ ($c_{a}>c_{b}$), which fits with the idea that the process is possible since the particle $a$ has a greater phase velocity than the particle $b$.
@@
Before doing so, we can study the relation between $\mathbb{A}$ and $\mathbb{B}$ when the emitted particle $b$ has arbitrarily low momentum, $k\to 0$, or carries off all of the momentum of the particle $a$, $k\to p$. In the former case, it is easy to see from Eq. (15) that Cherenkov emission is possible if
@@
The threshold condition is recovered setting an equal sign in Eqs. (17) and (18).
@@
It is easy to check that, for the physical momentum $k\in[0,p]$, $g^{\prime}(k)>0$ and $g(0)$, $g(p)>0$; hence $g(k)>0$. The remaining terms in Eq. (19) are positive, so we also have that $f^{\prime}(k)>0$; in other words, $f(k)$ is a monotonically increasing function. Therefore, if $f(0)<0$ and $f(p)>0$, the solution $k_{\text{th}}$ is unique. The condition $f(0)<0$ is satisfied when Eq. (17) does not hold, and $f(p)>0$ when Eq. (18) is fulfilled. The threshold momentum $k_{\text{th}}$ here establishes the minimum momentum that particle $b$ must have in order to have Cherenkov emission with momentum $k\in[k_{\text{th}},p]$, and it is permitted as long as $k_{\text{th}}<p$. The limit case $k_{\text{th}}=p$ occurs setting an equal sign in Eq. (18).
The proof is similar when $\mathbb{A}<0$ and $\mathbb{B}<0$. In this case one can check that $f(k)$ is a monotonically decreasing function, and so there exists a unique solution $k_{\text{th}}$ when $f(0)>0$ and $f(p)<0$. The first condition is met when Eq. (17) is satisfied, and the second one when Eq. (18) is not fulfilled. Hence, $k_{\text{th}}$ establishes the maximum momentum that particle $b$ can have up to where the Cherenkov effect is allowed, and thus $k\in[0,k_{\text{th}}]$. The process is permitted when $k_{\text{th}}>0$, and the limit case $k_{\text{th}}=0$ is given setting an equal sign in Eq. (17).
@@
Table 1: In this table we show the threshold momentum $k_{\text{th}}$ for the particle $b$ that satisfies the condition $f(k)=0$ together with the conditions that $\mathbb{A}$ and $\mathbb{B}$ should satisfy for the process to be allowed. The $\pm$ signs in the quadratic solutions depend on the signs of $\mathbb{A}$ and $\mathbb{B}$, and these are chosen so that $k_{\text{th}}$ is positive for those LIVPs that fulfill either the condition (17) or (18). We emphasize again that $k_{\text{th}}$ establishes the minimum momentum of the particle $b$ when $\mathbb{A},\mathbb{B}>0$, and a maximum momentum when $\mathbb{A},\mathbb{B}<0$.
The threshold condition $\Theta_{c}=0$ does not have a general solution $k_{\text{th}}$ for arbitrary $\alpha$ and $\beta$ since it is a polynomial Eq. of degree $\max\{\alpha+1,\beta+1\}$. Nevertheless, particular solutions for small values of $\alpha$ and $\beta$ can be derived, and further simplified in the massless limit $m\to 0$. For example, consider the case where $\alpha=1$ and $\beta=0$. One easily checks from Eq. (16) that the threshold momentum for $m=0$ is given by